nimoali:

Bobby and Chuck are playing in the back yard, busily filling a big bucket with mud. If they could fill the bucket together in 15 minutes and it would take Bobby 40 minutes to fill it alone, how long would it take Chuck to fill the bucket?

justjm:

-If it takes Bobby to do the job in 40 minutes, that means that in 1 minute, Bobby's completed 1/40th of the job -If it takes Bobby and Chuck to finish the job in 15 minutes, that means in 1 minute, 1/15th of the job is done together. -Hence, in 1 minute, the job completed for Chucks is 1/x You can form an equation \(\frac{1}{x} + \frac{1}{40} = \frac{1}{15}\) Now you can solve for x to find the time it takes for Chuck to finish the job.